OpenStudy (precal):

limit question as h approaches 0

OpenStudy (precal):

\[\frac{ e^(a+h) -e^a }{ h }\]

OpenStudy (precal):

that should be e to the (a+h) power

OpenStudy (anonymous):

^{a+h}

OpenStudy (anonymous):

use that latex.

OpenStudy (precal):

\[\frac{ e^{a+h}-e^a }{ h }\]

OpenStudy (anonymous):

\[\large\frac{e^{a+h}-e^n}{h}\]

OpenStudy (precal):

sorry I don't know latex but I use the the braces. Don't forget the second term is e^a

OpenStudy (precal):

I factored out e^a

OpenStudy (precal):

not sure what to do, I sub h=0 and I got e^a (0/0)

OpenStudy (anonymous):

\[\huge\lim_{h \rightarrow 0}~ \frac{e^{a+h}-e^{a} }{h}\]

OpenStudy (precal):

I was told the solution is e^a but I don't know how

OpenStudy (precal):

does 0/0 cancel out? but I thought 0/0 is undefined

OpenStudy (precal):

I have a Ti-89, I don' t know how to input a limit, it does do limits. The worksheet is suppose to be done with a graphing calculator but I don't know how to do this limit.

OpenStudy (anonymous):

I can't get farther than